OpenStudy (anonymous):

x: 3(x + 1) = −2(x − 1) − 4 someone help please???

OpenStudy (anonymous):

Distribute 3 to the brackets first and tell me what did you get ??

OpenStudy (anonymous):

x=-1

OpenStudy (anonymous):

Distribute means : \[a(b+c) = ab + ac\]

OpenStudy (anonymous):

3x + 3 = -2x - 2 ?

OpenStudy (anonymous):

Sorry, I thought you have just distributed that. Yeah you are right... Add 2x both the sides then..

OpenStudy (anonymous):

@gerryliyana let him or her try first...

OpenStudy (anonymous):

ohh i got it!! can someone help me with one last one? 3x − 5 = 2x + 6

OpenStudy (anonymous):

Subtract 2x both the sides..

OpenStudy (anonymous):

1x-5=x+6 ?

OpenStudy (anonymous):

Okay, you have to get the "x" to ONE side, so you have to minus one of them and than you have to do it to the other side of the = singe so.... #1- 3x + 5 = 2x + 6 -2x -2x ______________________________________… #2- x + 5 = 6 You want to get x by itself, so you -5 from both sides -5 -5 ______________________________________… #3- x = 1 Theres your answer!! Hope that helped!! =] =]

OpenStudy (anonymous):

|-2x | 3x − 5 = 2x + 6 |-2x|

OpenStudy (anonymous):

Mistake..

OpenStudy (anonymous):

\(3x - 2x - 5 = 2x - 2x + 6\)

OpenStudy (anonymous):

so the answer is one?

OpenStudy (anonymous):

It becomes : \(x - 5 = 6\) Now add 5 both the sides you will get x...

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@ANAS23 now i get it! thanks :)

OpenStudy (anonymous):

@ANAS23 answer is not 1...

OpenStudy (anonymous):

Original question says -5 but you have taken +5 instead...

OpenStudy (anonymous):

\(x-5 = 6\) Add 5 both the sides to find x...

OpenStudy (anonymous):

and your answer waterineyes

OpenStudy (anonymous):

and your answer @waterineyes

OpenStudy (anonymous):

What is in the answer @ANAS23 ? Just understand the solution..

OpenStudy (anonymous):

@ANAS23 what will you get on adding 5 to both the sides of : \(x-5 = 6\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

11

OpenStudy (anonymous):

Yes then not 1 but double 1 is the answer i.e. 11..